OpenStudy (anonymous):

2/x-2 + 4/x+1 = 3???

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {a \over c} + {b \over d} = {ad + bc \over cd } }\)

OpenStudy (mertsj):

Multiply all terms by the common denominator (x-2)(x+1)\[(x-2)(x+1)\times\frac{2}{x-2}+(x-2)(x+1)\times\frac{4}{x+1}=3(x-2)(x+1)\]

OpenStudy (mertsj):

\[2(x+1)+4(x-2)=3(x^2-x-2)\]

OpenStudy (mertsj):

\[2x+2+4x-8=3x^2-3x-6\]

OpenStudy (mertsj):

\[0=3x^2-9x\]

OpenStudy (mertsj):

\[3x(x-3)=0\]

OpenStudy (mertsj):

\[x=0, x=3\]

OpenStudy (anonymous):

2-2/x=4/x+1=3 2-6/x+1=3 6/x-1=3 6/x=4 x=6/4